# A Unified Framework for Analyzing Meta-algorithms in Online Convex Optimization

### v3

@@ G Proof of Theorem 8
where ${\mathbf{v}}$ is a random vector, taking its values uniformly from ${\mathbb{S}}^{1}\cap{\mathcal{L}}_{0}={\mathbb{S}}^{1}\cap(\operatorname{aff}({\mathcal{K}})-{\mathbf{z}})$, for any ${\mathbf{z}}\in{\mathcal{K}}$ and $k=\operatorname{dim}({\mathcal{L}}_{0})$. Since ${\mathcal{Q}}_{t}$ is a stochastic value oracle for $f_{t}$, according to Lemma 5 in (?), $\hat{{\mathcal{Q}}}_{t}({\mathbf{x}})$ is an unbiased estimator of $\nabla\hat{f}_{t}({\mathbf{x}})$. Moreover, we have
$\|\frac{k}{2\delta}\left(f_{t}({\mathbf{x}}+\delta{\mathbf{v}})-f_{t}({\mathbf{x}}-\delta{\mathbf{v}})\right){\mathbf{v}}\|\leq\frac{k}{2\delta}M_{1}\|({\mathbf{x}}+\delta{\mathbf{v}})-({\mathbf{x}}-\delta{\mathbf{v}})\|\leq kM_{1}.$
Hence we have $\hat{B}\in\operatorname{Adv}_{1}^{\text{o}}(\hat{{\mathbf{F}}},kM_{1})$. Using Equation 12 and the definition of the Meta-algorithm 4, we see that the responses of the queries are the same between the game $({\mathcal{A}},\hat{{\mathcal{B}}})$ and $({\mathcal{A}}^{\prime},{\mathcal{B}})$. It follows that the sequence of actions $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{T})$ in $({\mathcal{A}},\hat{{\mathcal{B}}})$ corresponds to the same sequence of actions in $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
@@
According to Lemma 3 in (?), we have $|\hat{f_{t}}({\mathbf{x}}_{t})-f_{t}({\mathbf{x}}_{t})|\leq\delta M_{1}$. On the other hand, we have
$\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$ $=\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$
$\leq\delta M_{1}T+\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{u}}}_{t})$ (Lemma 3 in (?))
$=\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}\left(\left(1-\frac{\delta}{r}\right){\mathbf{u}}_{t}+\frac{\delta}{r}{\mathbf{c}}\right)$ (Definition of $\hat{{\mathcal{U}}}$)
$=\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{K}}}\sum_{t=a}^{b}f_{t}\left({\mathbf{u}}_{t}+\frac{\delta}{r}({\mathbf{c}}-{\mathbf{x}})\right)$
$\leq\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(f_{t}({\mathbf{u}}_{t})+\frac{2\delta M_{1}D}{r}\right)$ (Lipschitz)

### v4

@@ Appendix G Proof of Theorem 14
where $\mathbf{v}$} is a random vector, taking its values uniformly from ${\mathbb{S}}^{1}\cap{\mathcal{L}}_{0}={\mathbb{S}}^{1}\cap(\operatorname{aff}({\mathcal{K}})-{\mathbf{z}})$, for any ${\mathbf{z}}\in{\mathcal{K}}$ and $k=\operatorname{dim}({\mathcal{L}}_{0})$. Since ${\mathcal{Q}}_{t}$ is a stochastic value oracle for $f_{t}$, according to Lemma 5 in pedramfar23_unified_approac_maxim_contin_dr_funct, $\hat{{\mathcal{Q}}}_{t}({\mathbf{x}})$ is an unbiased estimator of $\nabla\hat{f}_{t}({\mathbf{x}})$. Moreover, we have
$\left\|\frac{k}{2\delta}\left(f_{t}({\mathbf{x}}+\delta{\mathbf{v}})-f_{t}({\mathbf{x}}-\delta{\mathbf{v}})\right){\mathbf{v}}\right\|\leq\frac{k}{2\delta}M_{1}\|({\mathbf{x}}+\delta{\mathbf{v}})-({\mathbf{x}}-\delta{\mathbf{v}})\|\leq kM_{1}.$
[…]
@@
According to Lemma 3 in pedramfar23_unified_approac_maxim_contin_dr_funct, we have $|\hat{f_{t}}({\mathbf{x}}_{t})-f_{t}({\mathbf{x}}_{t})|\leq\delta M_{1}$. On the other hand, we have
$\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$ $=\min_{\hat{{\mathbf{w}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{w}}}_{t})$
$\leq\delta M_{1}T+\min_{\hat{{\mathbf{w}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{w}}}_{t})$ (Lemma 3 in pedramfar23_unified_approac_maxim_contin_dr_funct)
$=\delta M_{1}T+\min_{{\mathbf{w}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}\left(\left(1-\frac{\delta}{r}\right){\mathbf{w}}_{t}+\frac{\delta}{r}{\mathbf{c}}\right)$ (Definition of $\hat{{\mathcal{U}}}$)
$=\delta M_{1}T+\min_{{\mathbf{w}}\in{\mathcal{K}}}\sum_{t=a}^{b}f_{t}\left({\mathbf{w}}_{t}+\frac{\delta}{r}({\mathbf{c}}-{\mathbf{w}}_{t})\right)$
$\leq\delta M_{1}T+\min_{{\mathbf{w}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(f_{t}({\mathbf{w}}_{t})+\frac{2\delta M_{1}D}{r}\right)$ (Lipschitz)